Constants Part A 550 V is applied to a 2900-pF capacitor. How much energy is stored? Express your answer using three significant figures and include the a ppropriate units. PEValue Units Submit est Answer Provide Feedback Next>
Answer #1

given

V = 550 V
C = 2900 pF

= 2900*10^-12 F

Potential energy stored, PE = (1/2)*C*V^2

= (1/2)*2900*10^-12*550^2

= 4.39*10^-4 J